Grade 12th pass10 grade maths

Ex 6.1 Class 8 Maths Question 8. (i) Express 49 as the sum of 7 odd numbers. (ii) Express 121 as the sum of 11 odd numbers

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Pawan Prajapati
5 Years ago
Solution: (i) 49 = 1 + 3 + 5 + 7 + 9 + 11 + 13 (n = 7) (ii) 121 = 1 + 3 + 5 + 7 + 9 + 11 + 13 + 15 + 17 + 19 + 21 (n = 11)